# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

### v1

@@ V Correlator matrix and surrogate environment
and $\hat{U}_{E}(t)=\exp(-i\hat{H}_{E}^{bright}t/\hbar)$ is the free evolution operator of the bright bosonic modes of the electromagnetic environment. Using this property, it is possible to introduce an equivalent surrogate environment with only $N$ bosonic modes that reproduces the same reduced dynamics of the quantum emitters.
@@
$C_{ij}(t)=C^{(M)}_{ij}(t)+C^{(S)}_{ij}(t).$ (59)
This is the case, for instance, when both reservoirs are initially in vacuum, thermal, or number states. In such situations, the influence of the electromagnetic environment on the emitter ensemble is fully characterized by the $N\times N$ correlator matrix $C(t)=[C_{ij}(t)]$.
V.1 Correlators for Thermal Initial States
We now consider scenarios in which: (a) the initial quantum state of the electromagnetic environment is a product state, that is, $\hat{\rho}_{E}(0)=\hat{\rho}_{E}^{(M)}(0)\otimes\hat{\rho}_{E}^{(S)}$, where $\hat{\rho}_{E}^{(M)}(0)$ and $\hat{\rho}_{E}^{(S)}(0)$ are the initial density operators of the medium-assisted reservoir and of the scattering-assisted reservoir, respectively; (b) the two reservoirs are initially in thermal quantum states with temperatures $T_{0}^{(M)}$ and $T_{0}^{(S)}$. Under these assumptions, the correlation function $C_{ij}^{(\alpha)}(t)$ with $\alpha=M,S$ becomes (see Appendix C for details):
$C_{ij}^{(\alpha)}(t)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{0}^{\infty}\mathrm{d}\omega[(1+n_{\omega}^{(\alpha)})\mathcal{J}_{ij}^{(\alpha)}(\omega)e^{-i\omega t}+\\ n_{\omega}^{(\alpha)}\mathcal{J}_{ij}^{(\alpha)*}(\omega)e^{+i\omega t}]$ (60)
where
$n_{\omega}^{(\alpha)}=\frac{1}{e^{\beta_{\alpha}\hbar\omega}-1},$ (61)
and $\beta_{\alpha}=1/(k_{B}T_{0}^{(\alpha)})$. It is convenient to rewrite 60 as
$C^{(\alpha)}_{ij}(t)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{-\infty}^{+\infty}\mathrm{d}\omega\,\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})e^{-i\omega t},$ (62)
@@ V.1 Correlators for Thermal Initial States
The function $\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})$ is continuous at $\omega=0$ because the spectral density matrices go to zero for $\omega\rightarrow 0$ at least as $\omega^{2}$. Eventually, the correlation function $C_{ij}(t)$ can be expressed as
$C_{ij}(t)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{-\infty}^{+\infty}d\omega\mathrm{J}^{eff}_{ij}(\omega;\beta_{M},\beta_{S})e^{-i\omega t},$ (64)
@@
where now
$n_{\omega}=\frac{1}{e^{\beta\hbar\omega}-1},$ (67)
and $\beta=1/(k_{B}T_{0})$. In the zero-temperature limit, expression 66 reduces to
$\quad\;\mathrm{J}^{\text{eff}}_{ij}(\omega;\infty)=\begin{cases}\frac{1}{2\pi}\Gamma_{ij}(\omega)&\text{for}\,\omega\geq 0,\\ 0&\text{for}\,\omega\leq 0.\end{cases}$ (68)
@@ V.2 Surrogate Environment Initially in Vacuum Quantum State
Let $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}(\omega;\beta_{S})$ denote the $N\times N$ complex matrices with elements $\mathrm{J}^{(M)}_{ij}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}_{ij}(\omega;\beta_{S})$, respectively. We call these matrices temperature-dependent spectral density matrices of the medium and scattering-assisted reservoirs, respectively. These matrices are defined for $\infty<\omega<+\infty$, Hermitian, and positive definite. We now introduce the effective spectral density matrix
@@
which characterizes the overall electromagnetic environment. This matrix is also defined for $\infty<\omega<+\infty$, complex, Hermitian, and positive definite.
@@
$\hat{H}^{sur}=\hat{H}_{A}+\hat{H}_{E}^{sur}+\hat{H}_{I}^{sur}$ (70)
where $\hat{H}_{E}^{sur}$ is the bare Hamiltonian of a surrogate environment consisting of $N$ bosonic reservoirs, with positive and negative frequencies, initially in the vacuum state,
$\hat{H}_{E}^{sur}=\sum_{j=1}^{N}\int_{-\infty}^{+\infty}d\omega\hbar\omega\hat{a}_{j}^{\dagger}(\omega)\hat{a}_{j}(\omega);$ (71)
here $\hat{a}_{j}(\omega)$ and $\hat{a}_{j}^{\dagger}(\omega)$, with $j=1,N$, are the annihilation and creation operators of the bosonic modes of the surrogate environment. The interaction Hamiltonian $\hat{H}_{I}^{sur}$ between the quantum emitters and the surrogate environment is given by
$\hat{H}_{I}^{surr}=-\sum_{i,j=1}^{N}\hat{\mu}_{i}\int_{-\infty}^{+\infty}d\omega[G^{\text{eff}}_{ij}(\omega)\hat{a}_{j}(\omega)+h.c.],$ (72)
@@
${G}^{\text{eff}}(\omega)=\hbar[D^{-1}{J}^{\text{eff}}(\omega;\beta_{M},\beta_{M})D^{-1}]^{1/2}.$ (73)
When the medium-assisted reservoir and the scattering-assisted reservoir are initially in the vacuum state, the effective spectral density $\mathrm{J}_{ij}^{\text{eff}}(\omega)$ is given by 68. In the literature relevant to multiple quantum emitters that interact with dielectric objects based on the Langevin noise formalism, this expression is used [8]. However, as already pointed out in [1], [2] for a single quantum emitter, the conditions under which 68 remains valid are not always clearly stated. In fact, when the two reservoirs are initially in thermal quantum states, the expression 68 is not more valid. If the temperatures of the two reservoirs are the same, the elements of the effective spectral density matrix are given by 66. When the initial temperatures are different, the elements of the effective spectral density matrix are given by 65.
@@
Figure 2: Two equal quantum emitters with dipole moment $\mu$ interact interact with a nanostructure composed of two rods and a disk. The first rod has length $3a$, width $2a$, and height $a$; the second rod has length $4a$, width $2a$, and height $a$; the disk has radius $a$ and height $a$. All three particles are made of a Drude material with plasma frequency $\omega_{p}$ and damping rate $\nu$ such as $\nu/\omega_{p}=0.01$. The two quantum emitters are placed at the midpoint of the gaps between adjacent particles. Normalized spectral densities $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ versus $\omega/\omega_{p}$ for $i,j=1,2$ and $k_{p}\,a=1$ where $k_{p}=\omega_{p}/c$. The characteristic spectral density $\mathcal{J}_{0}$ is given by $\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2}$. The spectral density matrices are Hermitian.
@@ VI Application to a system with two quantum emitters
In order to showcase the effects of an electromagnetic environment consisting of dispersive dielectric particles and vacuum on a system made of multiple quantum emitters, we examine the minimal setting of two quantum emitters, labeled 1 and 2. We consider two representative electromagnetic environments: a spherical particle in vacuum, as in [2], and a nanostructure composed of two rods and a disk in vacuum (see the insets of Figs. 1 and 2, respectively). The two quantum emitters are placed at the positions indicated by the red dots in the insets. All particles are homogeneous and their dielectric permittivity is described by the Drude model: $\varepsilon_{\omega}=[1-\omega_{p}^{2}/(\omega^{2}+i\nu\omega)]$ with plasma frequency $\omega_{p}$ and relaxation frequency $\nu$.
Within the modified Langevin noise formalism and the emitter-centered description, the electromagnetic environment is modeled by two continuous bosonic baths, medium (M)- and scattering (S)- assisted reservoirs, characterized by the spectral density matrices $\mathcal{J}^{(M)}(\omega)$ and $\mathcal{J}^{(S)}(\omega)$. The two reservoirs are initially in thermal quantum states at inverse temperatures $\beta_{\text{M}}$ and $\beta_{\text{S}}$. The two emitters interact through their dipole moments with the continuous sets of bosonic excitations of the electromagnetic environment.
@@
$\hat{H}_{A}=\frac{\hbar\Omega_{1}}{2}\sigma_{z}^{(1)}+\frac{\hbar\Omega_{2}}{2}\sigma_{z}^{(2)},$ (74)
and the transition dipole moment operators are $\hat{\mu}_{1}=\mu_{1}\hat{\sigma}^{(1)}_{x}$, $\hat{\mu}_{2}=\mu_{2}\hat{\sigma}^{(2)}_{x}$; here $\sigma_{z}^{(i)}=|e^{(i)}\rangle\langle e^{(i)}|-|g^{(i)}\rangle\langle g^{(i)}|$ and $\sigma_{x}^{(i)}=|g^{(i)}\rangle\langle e^{(i)}|+|e^{(i)}\rangle\langle g^{(i)}|$ where $|g^{(i)}\rangle$ and $|e^{(i)}\rangle$ are the ground and excited states of the $i$-th quantum emitter. In the numerical examples below, we considered equal dipole moments $\mu_{1}=\mu_{2}\equiv\mu$.
As shown in Sec. V, leveraging the temperature-dependent spectral densities $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}(\omega;\beta_{S})$, the degrees of freedom of the medium-assisted and scattering-assisted baths are reduced to those of two distinct sets of surrogate continuous bosonic reservoirs initially in the vacuum quantum state and with an effective spectral density matrix $\mathrm{J}^{\text{eff}}(\omega;\beta_{M},\beta_{S})$. Furthermore, each emitter is coupled to both continuous reservoirs through direct coupling $G^{\text{eff}}_{ii}(\omega)$ and cross-coupling amplitudes $G^{\text{eff}}_{ij}(\omega)$. The coupling of two emitters with a common reservoir has been known to generate quantum correlated states of the composite system [25], similar to what has been studied in the field of cQED [26]. However, any real electromagnetic environment with linear constitutive relations is characterized by a number of reservoirs equal to the number of emitters: each emitter is simultaneously coupled to all of them. We study the non-equilibrium dynamics of the emitters for a set of physically relevant initial quantum states, including pure product states and maximally entangled Bell states.
We investigate the reduced dynamics of the two quantum emitters using the Matrix Product State approach, similarly to what is done with the TEDOPA method [23], to simulate the time evolution of the whole emitter + surrogate environment state governed by the effective Hamiltonian defined by Eqs. 70-74. Such methods have been shown to reproduce exact results and to provide corrections to more recent numerical techniques based on collisional models [27].
@@ VI.1 Spectral Densities
$\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2}$ (75)
@@
Figure 1 shows $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ for a Drude sphere with radius $a$ as a function of the normalized frequency $\omega/\omega_{p}$, assuming size parameter $k_{p}\,a=1$ and normalized damping rate $\nu/\omega_{p}=0.01$. The two emitters are positioned as in the inset, each at a distance $d$ from the center of the sphere, with $d/a=1.75$. In the low–frequency region around the dipolar plasmon resonance ($\omega/\omega_{p}\simeq 0.54$), the amplitude of the elements $\mathcal{J}^{(S)}_{ij}$ is greater than or equal to $\mathcal{J}^{(M)}_{ij}$. At higher frequencies, near higher-order plasmonic resonances, the amplitude of $\mathcal{J}^{(M)}_{ij}$ becomes significantly larger than that of $\mathcal{J}^{(S)}_{ij}$. For a comprehensive analysis of the role of the scattering modes in shaping the overall spectral density, see Ref. [2]. The bandwidths of the plasmonic modes are limited by a combination of radiative and material losses, as detailed in [2].
Figure 2 shows the corresponding normalized spectral densities for a nanostructure composed of two rods and a disk. The first rod has length $3a$, width $2a$, and height $a$; the second rod has length $4a$, width $2a$, and height $a$; the disk has radius $a$ and height $a$. The emitters are placed at the midpoint of the gaps between adjacent particles. All materials follow the same Drude model with $k_{p}\,a=1$ and a damping rate $\nu/\omega_{p}=0.01$. Also in this case the scattering-assisted spectral densities are typically greater than medium-assisted at low frequencies, while at low frequencies the material-assisted contribution dominates and is typically characterized by sharper peaks associated to higher multipolar scattering orders.

### v2

@@ V Correlator matrices and surrogate environment
and $\hat{U}_{E}(t)=\exp(-i\hat{H}_{E}^{\text{bright}}t/\hbar)$ is the free evolution operator of the bright bosonic modes of the electromagnetic environment. Using this property, it is possible to introduce an equivalent surrogate environment with only $N$ bosonic modes that reproduces the same reduced dynamics of the quantum emitters.
@@
$C_{ij}=C^{(M)}_{ij}+C^{(S)}_{ij}$. (59)
This is the case, for instance, when both reservoirs are initially in vacuum, thermal, or undisplaced squeezed states. In such situations, the influence of the electromagnetic environment on the emitter ensemble is fully characterized by the $N\times N$ correlator matrix $C(t+\tau;t)=[C_{ij}(t+\tau;t)]$.
V.1 Reservoirs Initially in Thermal Quantum States
We now consider scenarios in which the two reservoirs are initially in thermal quantum states at temperatures $T_{0}^{(M)}$ and $T_{0}^{(S)}$. The corresponding correlators are then stationary, i.e., $C_{ij}^{(\alpha)}(\tau)\equiv C_{ij}^{(\alpha)}(t+\tau;t)$ for $\alpha=M,S$, and read (see Appendix C):
$C_{ij}^{(\alpha)}(\tau)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{0}^{\infty}\mathrm{d}\omega[(1+n_{\omega}^{(\alpha)})\mathcal{J}_{ij}^{(\alpha)}(\omega)e^{-i\omega\tau}+\\ n_{\omega}^{(\alpha)}\mathcal{J}_{ij}^{(\alpha)*}(\omega)e^{+i\omega\tau}],$ (60)
where $n_{\omega}^{(\alpha)}=({e^{\beta_{\alpha}\hbar\omega}-1})^{-1}$, and $\beta_{\alpha}=1/(k_{B}T_{0}^{(\alpha)})$. It is convenient to rewrite 60 as
$C^{(\alpha)}_{ij}(\tau)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{-\infty}^{+\infty}\mathrm{d}\omega\,\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})e^{-i\omega\tau},$ (61)
@@ V.1 Reservoirs Initially in Thermal Quantum States
The function $\mathrm{J}^{(\alpha)}_{ij}(\omega;\beta_{\alpha})$ is continuous at $\omega=0$ because the spectral density matrices go to zero for $\omega\rightarrow 0$ at least as $\omega^{2}$. Consequently, the correlation function $C_{ij}(\tau)$ can be expressed as
$C_{ij}(\tau)=\frac{\hbar^{2}}{\mu_{i}\mu_{j}}\int_{-\infty}^{+\infty}d\omega\mathrm{J}^{\text{eff}}_{ij}(\omega;\beta_{M},\beta_{S})e^{-i\omega\tau},$ (63)
@@
where now $n_{\omega}=(e^{\beta\hbar\omega}-1)^{-1},$ and $\beta=1/(k_{B}T_{0})$. In the zero-temperature limit, expression 65 reduces to
$\quad\;\mathrm{J}^{\text{eff}}_{ij}(\omega;\beta\rightarrow\infty)=\begin{cases}\frac{1}{2\pi}\Gamma_{ij}(\omega)&\text{for}\,\omega\geq 0,\\ 0&\text{for}\,\omega\leq 0.\end{cases}$ (66)
@@ V.2 Surrogate Environment Initially in Vacuum Quantum State
Let $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}(\omega;\beta_{S})$ denote the $N\times N$ complex matrices with elements $\mathrm{J}^{(M)}_{ij}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}_{ij}(\omega;\beta_{S})$, respectively. We call these matrices temperature-dependent spectral density matrices of the medium and scattering-assisted reservoirs, respectively. These matrices are defined for $\infty<\omega<+\infty$, they are Hermitian and positive definite. We now introduce the effective spectral density matrix
@@
which characterizes the overall electromagnetic environment. This matrix is also defined for $\infty<\omega<+\infty$, Hermitian, and positive definite.
@@
$\hat{H}^{\text{sur}}=\hat{H}_{A}+\hat{H}_{E}^{\text{sur}}+\hat{H}_{I}^{\text{sur}},$ (68)
where $\hat{H}_{E}^{\text{sur}}$ is the bare Hamiltonian of a surrogate environment consisting of $N$ bosonic reservoirs, with positive and negative frequencies, initially in the vacuum state,
$\hat{H}_{E}^{\text{sur}}=\sum_{j=1}^{N}\int_{-\infty}^{+\infty}d\omega\hbar\omega\hat{a}_{j}^{\dagger}(\omega)\hat{a}_{j}(\omega);$ (69)
here $\hat{a}_{j}(\omega)$ and $\hat{a}_{j}^{\dagger}(\omega)$, with $j=1,\ldots,N$, are the annihilation and creation operators of the bosonic modes of the surrogate environment. The interaction Hamiltonian $\hat{H}_{I}^{\text{sur}}$ between the quantum emitters and the surrogate environment is given by
$\hat{H}_{I}^{\text{sur}}=-\sum_{i,j=1}^{N}\hat{\mu}_{i}\int_{-\infty}^{+\infty}d\omega[G^{\text{eff}}_{ij}(\omega)\hat{a}_{j}(\omega)+h.c.],$ (70)
@@
${G}^{\text{eff}}(\omega)=\hbar[P^{-1}\mathrm{J}^{\text{eff}}(\omega;\beta_{M},\beta_{S})P^{-1}]^{1/2}.$ (71)
When the medium-assisted reservoir and the scattering-assisted reservoir are initially in the vacuum state, the effective spectral density $\mathrm{J}_{ij}^{\text{eff}}(\omega)$ is given by 66. In the literature on multiple quantum emitters interacting with dielectric objects based on the macroscopic quantum electrodynamic approach using the Langevin noise formalism, this expression is widely used (e.g. [8], [31]). However, as already pointed out in [1] and [2] for a single quantum emitter, the conditions under which 66 remains valid are not always clearly stated. In fact, when the two reservoirs are initially in thermal quantum states at non-zero temperature, the expression 66 is no longer valid. If the temperatures of the two reservoirs are the same, the elements of the effective spectral density matrix are given by 65. When the initial temperatures are different, the elements of the effective spectral density matrix are given by 64.
@@
Figure 2: Two equal quantum emitters with dipole moment $\mu$ interact with a nanostructure composed of two rods and a disk. The first rod has length $3a$, width $2a$, and height $a$; the second rod has length $4a$, width $2a$, and height $a$; the disk has radius $a$ and height $a$. All three particles are made of a Drude material with plasma frequency $\omega_{p}$ and damping rate $\nu$ such as $\nu/\omega_{p}=0.01$. The two quantum emitters are placed at the midpoint of the gaps between adjacent particles. Normalized spectral densities $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ versus $\omega/\omega_{p}$ for $i,j=1,2$ and $k_{p}\,a=1$ where $k_{p}=\omega_{p}/c$. The characteristic spectral density $\mathcal{J}_{0}$ is given by $\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2}$. The spectral density matrices are Hermitian.
@@ VI Application to a system with two quantum emitters
We illustrate the impact of a structured electromagnetic environment, formed by dispersive dielectric particles embedded in vacuum, on the dynamics of an emitter ensemble by focusing on the simplest nontrivial case: two quantum emitters, labeled 1 and 2. We consider two representative nanophotonics settings: a spherical particle in vacuum, as in [2], and a nanostructure composed of two rods and a disk in vacuum (see the insets of Figs. 1 and 2, respectively). The two quantum emitters are placed at the positions indicated by the red dots in the insets. All particles are homogeneous and their dielectric permittivity is described by the Drude model: $\varepsilon_{\omega}=[1-\omega_{p}^{2}/(\omega^{2}+i\nu\omega)]$ with plasma frequency $\omega_{p}$ and relaxation frequency $\nu$.
Within the modified Langevin noise formalism and the emitter-centered description, the electromagnetic environment is modeled by two continuous bosonic baths, medium-assisted (M) and scattering-assisted (S) reservoirs, characterized by the spectral density matrices $\mathcal{J}^{(M)}(\omega)$ and $\mathcal{J}^{(S)}(\omega)$. The two reservoirs are initially in thermal quantum states at inverse temperatures $\beta_{\text{M}}$ and $\beta_{\text{S}}$. The two emitters interact through their dipole moments with the continuous sets of bosonic excitations of the electromagnetic environment.
@@
$\hat{H}_{A}=\frac{\hbar\Omega_{1}}{2}\hat{\sigma}_{z}^{(1)}+\frac{\hbar\Omega_{2}}{2}\hat{\sigma}_{z}^{(2)},$ (72)
[…]
As shown in Sec. V, leveraging the temperature-dependent spectral densities $\mathrm{J}^{(M)}(\omega;\beta_{M})$ and $\mathrm{J}^{(S)}(\omega;\beta_{S})$, the degrees of freedom of the medium-assisted and scattering-assisted baths are reduced to those of two distinct sets of surrogate continuous bosonic reservoirs initially in the vacuum quantum state and with an effective spectral density matrix $\mathrm{J}^{\text{eff}}(\omega;\beta_{M},\beta_{S})$. Furthermore, each emitter is coupled to both continuous reservoirs through direct coupling $G^{\text{eff}}_{ii}(\omega)$ and cross-coupling amplitudes $G^{\text{eff}}_{ij}(\omega)$. The coupling of two emitters with a common reservoir has been known to generate quantum correlated states of the composite system [32], similar to what has been studied in the field of cQED [33]. More generally, for linear electromagnetic environments the effective description involves a number of reservoirs equal to the number of emitters, and each emitter is simultaneously coupled to all of them. We study the non-equilibrium dynamics of the emitters for a set of physically relevant initial quantum states, including pure product states and maximally entangled Bell states.
We investigate the reduced dynamics of the two quantum emitters using the Matrix Product State approach, similarly to what is done with the TEDOPA method [29], to simulate the time evolution of the whole emitter + surrogate environment state governed by the effective Hamiltonian defined by Eqs. 68-72. Such methods have been shown to reproduce exact results and to provide corrections to more recent numerical techniques based on collisional models [34].
@@ VI.1 Spectral Densities
$\mathcal{J}_{0}=\frac{1}{6\pi^{2}}\frac{k_{p}}{\hbar\,\varepsilon_{0}}\left({\mu k_{p}}\right)^{2},$ (73)
@@
Figure 1 shows $\mathcal{J}^{(S)}_{ij}/\mathcal{J}_{0}$ and $\mathcal{J}^{(M)}_{ij}/\mathcal{J}_{0}$ for a Drude sphere with radius $a$ as a function of the normalized frequency $\omega/\omega_{p}$, assuming size parameter $k_{p}\,a=1$ and normalized damping rate $\nu/\omega_{p}=0.01$. The two emitters are positioned as in the inset, each at a distance $d$ from the center of the sphere, with $d/a=1.75$. In the low–frequency region around the dipolar plasmon resonance ($\omega/\omega_{p}\simeq 0.54$), the amplitude of the elements $\mathcal{J}^{(S)}_{ij}$ is comparable or larger than $\mathcal{J}^{(M)}_{ij}$. At higher frequencies, near higher-order plasmonic resonances, the amplitude of $\mathcal{J}^{(M)}_{ij}$ becomes significantly larger than that of $\mathcal{J}^{(S)}_{ij}$. For a comprehensive analysis of the role of the scattering modes in shaping the overall spectral density, see Ref. [2]. The bandwidths of the plasmonic modes are limited by a combination of radiative and material losses, as detailed in [2].
Figure 2 shows the corresponding normalized spectral densities for a nanostructure composed of two rods and a disk. The first rod has length $3a$, width $2a$, and height $a$; the second rod has length $4a$, width $2a$, and height $a$; the disk has radius $a$ and height $a$. The emitters are placed at the midpoint of the gaps between adjacent particles. All materials follow the same Drude model with $k_{p}\,a=1$ and a damping rate $\nu/\omega_{p}=0.01$. Also in this case, the scattering-assisted spectral densities are typically larger than medium-assisted at low frequencies, whereas at higher frequencies the medium-assisted contribution dominates and is typically characterized by sharper peaks associated with higher multipolar scattering orders.